OpenStudy (anonymous):

In a paragraph of at least three sentences, explain how the graphs of the two equations in your group are the same and how they are different. Be sure to use complete sentences. Equations: y = -3/2 + 4 y = 3/2 + 4

OpenStudy (anonymous):

They are both lines. y = mx + b <== slope intercept form They have different slopes. (m1 = -3/2)(m2=3/2) They have the same y-intercepts (b = +4)

OpenStudy (anonymous):

Ah, I thought of that I just can't imagine that they would have you make three sentences out of it...

OpenStudy (anonymous):

yeah those are some of the quick ideas i thought of. BTW they are not the only ones, there are other ideas too. but you kinda get the idea dont you? Oh so those, just make them into complete sentences. If you need any more assistance, just ask :)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw :)